OpenStudy (anonymous):

Find the width of a rectangle with perimeter 20in and length 2in

OpenStudy (texaschic101):

Perimeter of a rectangle is : p = 2(w + h) or 2w + 2h so now you have : 20 = 2w + 2(2) can you solve for w now ?

OpenStudy (anonymous):

let l = length and b = width of rectangle, hence 2(l+b) = perimeter of rectangle i.e. 2(2+b) =20 i.e. 4+2b =20 i.e. 2b =20-4 i.e. 2b =16 i.e. b=16/2 i.e. b=8 Thus the width of the rectangle is 8 in

OpenStudy (texaschic101):

well..ok...there is your answer

OpenStudy (anonymous):

thank you